OpenStudy (anonymous):

A roller coaster follows the path show below with no friction until it reaches point A: a) When the roller coaster reaches the top of the 20m loop, what is the magnitude AND direction of the force exerted by the track on the roller coaster? b) When the roller coaster reaches the point a it simultaneously encounters friction(uk = .2) and begins compressing a spring(k=2000 N/m) Write an equation to find the amount the spring is compressed, d.

OpenStudy (anonymous):

who really knows

OpenStudy (anonymous):

im posting the graph lol

OpenStudy (vishweshshrimali5):

Ok I think I can help you out.

OpenStudy (vishweshshrimali5):

First lets discuss about part (a).

OpenStudy (vishweshshrimali5):

Okay ?

OpenStudy (vishweshshrimali5):

On the top point A, there are only two forces : (1) Weight of the roller coaster acting downwards. (2) Reaction force exerted by the roller coaster on the track in upwards direction.

OpenStudy (vishweshshrimali5):

Now, since the net force on roller coaster in vertical direction is zero. So, N = weight = 800*g Now, this reaction force and the force exerted by track on roller coaster will form Newton' third law pair and hence will have same magnitude but opposite direction.

OpenStudy (vishweshshrimali5):

So, the answer for the part (a) would be, the required force (say F), F = N = 800*g = ? in downward direction.

OpenStudy (vishweshshrimali5):

@a25526598 can you tell me what will be the magnitude of F?